2. What is the total concentration of all ions (the sum of their individual concentrations) in aqueous solutions of the following?  

  • Hint: Each substance is an electrolyte that dissociates in water to produce ions. Use the link table of polyatomic ions if you do not recognize the polyatomic ion. For example, Na2SO4 -> 2Na1+ + SO42- when it dissolves, so a 1 M soln of Na2SO4 is 2M in Na1+ and 1M in SO42- and the total concentration of ions is 2 M + 1 M = 3M.
0.0500-M CaCl2 ......M
0.0840-M Na3PO4 ......M
0.0220-M K2Cr2O7 ......M
Answer #1

PART 1

Consider dissociation of CaCl 2 in water . CaCl 2 (aq) \rightarrow Ca 2+ (aq) + 2 Cl - (aq)

From reaction, 1 mol CaCl 2\equiv 1 mol Ca 2+\equiv 2 mol Cl -

\therefore [Ca 2+] =[CaCl 2] = 0.0500 M & [Cl -] = 2 [CaCl 2] = 2 ( 0.0500 M) = 0.100 M

Hence, Total concentration of ions is 0.0500 M + 0.100 M = 0.150 M

PART 2

Consider dissociation of Na3PO4 in water . Na3PO4 (aq) \rightarrow 3 Na + (aq) + PO4 3- (aq)

From reaction, 1 mol Na3PO4\equiv 3 mol Na +\equiv 1 mol PO4 3-

\therefore [Na +] = 3 [Na3PO4] = 3 (0.0840 M ) = 0.252 M & [ PO4 3- ] = [Na3PO4] = 0.0840 M

Hence, Total concentration of ions is 0.252 M + 0.0840 M =0.336 M

PART 3

Consider dissociation of K2Cr2O7 in water . K2Cr2O7 (aq) \rightarrow 2 K + (aq) + Cr2O7 2- (aq)

From reaction, 1 mol K2Cr2O7 \equiv 2 mol K +\equiv 1 mol Cr2O7 2-

\therefore [K +] = 2 [K2Cr2O7] = 2 ( 0.0220 M) = 0.0440 M & [ Cr2O7 2- ] = [K2Cr2O7] = 0.0220 M

Hence, Total concentration of ions is 0.0440 M + 0.0220 M =0.0660 M

ANSWER :

0.0500 M CaCl 2 0.150 M
0.0840 M Na3PO4 0.336 M
0.0220 M K2Cr2O7 0.0660 M